OpenStudy (anonymous):

sin2xcosx + cos2xsinx=-1/2

OpenStudy (badhi):

remember \(\sin(A+B) = \sin A\cos B+\cos A \sin B \)

OpenStudy (anonymous):

I need to find the solutions in the interval (0,2pi) for the equation

OpenStudy (badhi):

before finding the solution, did u use the given identity and simplified which leads to the solution?

OpenStudy (owlcoffee):

In order to solve this, we have to know some trigonometric identities: \[\cos2x=\cos^2x-\sin^2x=1-\sin^2x=2\cos^2x-1\] \[\sin2x=2senxcosx\] and use them to take all the double angles, and convert the whole equation into a single trigonometric function. or we could use, as Bahdi said, the sum of of angles. \[Sin(A+B)=sinAcosB+cosAsinB\] And, we can see, that the structure actually works, so, given: \[\sin2xcosx+\cos2xsinx=-\frac{ 1 }{ 2 }\] if we call A=2x and B=x, then: \[sinAcosB+cosAsinB=-\frac{ 1 }{ 2 }\] and, transforming the identity, we can see that: \[\sin(A+B)=-\frac{ 1 }{ 2 }\] but, we said that A=2x, and B=x So therefore: \[\sin(2x+x)=-\frac{ 1 }{ 2 }\] I'll let you continue from this point.